MOA-II Obsevation Log Monthly Summary 2006/10
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 4475-4713 | 239 | 1275 | 76420 |
LMC | 648-708 | 61 | 677 | 201700 |
SMC | 232-249 | 18 | 133 | 39000 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 318 | 2085 | 317120 (88h05m20s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
60 | 63 | 53 | 44 | 185 | 36 | 52 | 43 | 193 | 35 | 71 | 45 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
39 | 34 | 33 | 43 | 38 | 33 | 32 | 36 | 42 | 65 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
43 | 43 | 43 | 43 | 42 | 44 | 60 | 57 | 57 | 53 | 45 | 46 |
Field13 | Field14 | Field15 | Field16 |
49 | 52 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
17 | 16 | 16 | 16 | 14 | 14 | 14 | 13 | 13 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2006/10/1: DARK:D1585(60sec) and D1589(300sec) are OK. / night time = 34020(sec)
2006/10/2: No observations / night time = 33840(sec)
2006/10/3: no observation / night time = 33600(sec)
2006/10/4: D1597(60sec), D1599(300sec) are OK. We give up 40sec dark because edge of chip2 is always strange. Use D1597(60sec) for 40sec exposure science frames. / night time = 33420(sec)
2006/10/5: D1611(60sec), D1623(300sec) are OK / night time = 33240(sec)
2006/10/6: D1628(60sec), D1631(200sec) and D1636(300sec) are OK. / night time = 33000(sec)
2006/10/7: D1641(60sec), D1642(300sec) are OK. / night time = 32760(sec)
2006/10/8: / night time = 32580(sec)
2006/10/9: D1646(300sec) is OK. Anemometer was broken by strong wind. / night time = 32280(sec)
2006/10/10: D1650(60sec), D1651(300sec) are OK. / night time = 32100(sec)
2006/10/11: / night time = 31860(sec)
2006/10/12: D1655(60sec) is OK. / night time = 31680(sec)
2006/10/13: Dome damaged / night time = 31440(sec)
2006/10/14: / night time = 31200(sec)
2006/10/15: D1657(300sec) is OK / night time = 31020(sec)
2006/10/16: D1662(60sec), D1663(300sec) & D1664(15sec) are OK. Flats F1174 to F1209 are taken. / night time = 30780(sec)
2006/10/17: D1668(60sec) & D1670(300sec) are OK. / night time = 30480(sec)
2006/10/18: No observation. / night time = 30300(sec)
2006/10/19: No observation / night time = 30060(sec)
2006/10/20: D1678(60sec) is OK / night time = 29820(sec)
2006/10/21: use Dark D1680-300 and D1683-60. Test observation of Magellan bridge was done. / night time = 29580(sec)
2006/10/22: Use DARK, D1684-60 and D1686-300 / night time = 29340(sec)
2006/10/23: No obs / night time = 29160(sec)
2006/10/24: use D1687-60, D1690-300 / night time = 28920(sec)
2006/10/25: use D1695-60, No good DARK for 300sec was taken. T=-54deg at the end of obs. / night time = 28620(sec)
2006/10/26: Chunichi was defeated..., use D1699-60, D1704-300 / night time = 28380(sec)
2006/10/27: use D1706-60, D1708-300, MinElogation to the moon was changed to 5deg from 7deg. sky was less 10,000ADU. should be changed due to the moon age. / night time = 28140(sec)
2006/10/28: No obs. / night time = 27900(sec)
2006/10/29: No obs. Flat were taken. / night time = 27660(sec)
2006/10/30: use D1717-60, gave up DARK for 300sec after 5 trials. / night time = 27420(sec)
2006/10/31: use D1723-300. / night time = 27180(sec)
TOTAL NIGHT TIME: 951780(sec)
OPERATING RATE: 42.1 %
GB: 13.4 %
LMC: 24.0 %
SMC: 4.7 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month